Lane 1 swimmer is 1.7 second ahead of Lane 2 swimmer. Lane one swimmer is traveling at 1.4m/sec and Lane 2 is traveling at 1.5m/sec. How many seconds will it take Lane 2 to catch Lane 1?
Answer by mananth(16946)

Lane 1 swimmer is 1.7 seconds ahead
his speed is 1.4 m/s
so he is 2.38 meters
Catch up speed =
Difference between speed = 1.5-1.4 = 0.1
t=d/r
=2.38/0.1
= 23.8 seconds